OpenStudy (anonymous):

find set of parametric equations for the line which passes through the point (3,4,7) that is perpendicular to the line with parametric equations x=1+t, y=3-t, and z=5t

OpenStudy (anonymous):

That's almost identical to last problem

OpenStudy (anonymous):

where did the 2,-1,-1 come from

OpenStudy (anonymous):

from plugging in -1, -1 for b and c respectively

OpenStudy (anonymous):

Dot product of two vector are zero if they are perpendicular {1,-1,5}.{a,b,c}=0 a-b+5c=0 a=b-5c This is vector for our line {b-5c,b,c} x=3+(b-5c)t y=4+bt z=7+ct